1. What is the cost-of-carry model for pricing futures and forward contracts? Provide a “fair” futures price for each of the following assets:

a. S&P 500 Index: Contract June (3 months away); Current value of S&P 500 2450; 3-month Libor=2%.
Answer #1

Cost of Carry model implies that the Future and Forward Prices of assets will be greater than the Current spot price .

As per this model, Future or Forward Price of an asset =Current Spot price +Cost of financing and holding the asset

In case of commodities the cost of storing is included in the cost of holding the assets.

a. For S&P futures, the cost of interest is added to the spot price to arrive at the fair future price.

Current Value =2450

Libor =2%

Interest cost for 3 months =2450*0.02=$50

Fair Future Price  of the contract =2450+50=$2,500